# Fovea Stacking: Imaging with Dynamic Localized Aberration Correction

### v1

@@ 3.4.1. ROI tiling
Aberration correction across the field of view can be achieved by dividing the FoV into regions of interest (ROI) and correcting local aberrations independently. To optimize each wavefront for local correction within a ROI at distance $D$, $M$ points are sampled - typically on a grid - within the ROI. The optimization loss, $L_{ROI}$, is defined as the mean PSF radius across these sampled points:
@@ 3.4.2. Joint optimization
Although full-aberration correction can be achieved by tiling a grid of independently corrected ROIs, this approach does not efficiently exploit the information available across the entire image stack. Therefore, to jointly optimize aberration correction across the full FoV with a limited image budget $N$, we propose to optimize the PSF radius grid across the stack collectively.
We first sample a grid of $H\times W$ points over the entire FoV at a given depth. Ray-tracing is then performed for each of the different DPP wavefronts, and the corresponding PSF radius is calculated, resulting in a PSF radius grid stack $\{r_{n}(i,j)\in R^{H\times W}|1\leq n\leq N\}$. Each element $r_{n}(i,j)=r_{n}(P_{ij})$ represents the PSF radius of the sampled point $P_{ij}$, computed using the n-th DPP wavefront. The overall image quality after fusion depends on the minimal PSF radius at each location across the stack, i.e. $r_{min}(i,j)=\min_{n}r_{n}(i,j)$. Therefore, to jointly optimize for all DPPs, we minimize the PSF grid stacking loss .
(4) $L_{gs}=\frac{1}{H\times W}\sum_{i,j}{\min_{n}r_{n}(i,j)}$
Essentially, the grid stacking loss ensures that for each wavefront, only the area where it performs best across the entire stack will be considered as its active region for optimization. However, optimization may yield some degenerate patterns where does not outperform all other waveforms anywhere in the image, thus having no active region at all. This usually happens when number of image budgets increases. To alleviate it, we encourage these DPPs to optimize on the hard regions, using $r_{min}$ as a weight:
(5) $L_{hr}=$ $\sum_{n^{\prime}}\frac{1}{H\times W}\sum_{i,j}{r_{n^{\prime}}(i,j)r_{min}(i,j)}$
$n^{\prime}\in\{n|\sum_{i,j}\mathds{1}\left(n,n^{*}\right)=0\}$
where $\mathds{1}(\cdot)$ is the identity function, and $n^{*}(i,j)=argmin_{n}r_{n}(i,j)$ is the mask index for the sampled point, indicating which DPP wavefront provides the best imaging quality for it. The overall loss for joint optimization is a sum of the two losses:
@@
The joint optimization scheme is validated using the same simulated optical system, as shown in Fig. 5. In the experiment, $64\times 64$ grids are sampled across the full FoV to jointly optimize five wavefront patterns. The resulting optimized PSF grids exhibit compensatory patterns due to the gradients being masked differently by the index $n^{*}$ to optimize each DPP wavefront. As expected, the PSF grid displays radial symmetry, which the dynamic optimization scheme efficiently exploits to cover the entire FoV within a limited imaging budget.
@@ 3.4.3. Efficiency under limited imaging budget
The efficiency of using the imaging budget is analyized in Fig 6. Given a limited number of images, DPP patterns are optimized, and the average PSF radius of the composed PSF across the stack (rPSF) is used to measure overall imaging quality (blue). ROI tiling strategies requiring grid numbers of $2\times 2$, $3\times 3$, and $4\times 4$ are also included for a comparison (red). Joint optimization significantly improves image quality with as few as 3-5 images, whereas ROI tiling proves less efficient. Beyond 10 images, further improvements in image quality become marginal.
@@ 4. DPP Control Model
where $W\in R^{K}$ represents the Zernike coefficients, and $V^{2}\in R^{63}$ is the element-wise square of the electrode voltages. The matrix $A\in R^{K\times 63}$ is the influence matrix obtained through calibration. To calculate the voltage corresponding to a desired set of Zernike coefficients $W$, it is therefore modeled by solving a constrained optimization problem such that the voltage is within the physical limit $[0,V_{max}]$:
(8) $\min|W-AV^{2}|_{2}^{2}\quad s.t.0<V<V_{max}.$
However, we found that this linear control strategy fails to accurately estimate wavefront deformation, particularly at larger deformation scales. To address this limitation, we propose a neural network-based control strategy that captures the non-linear behavior of the device and improve accuracy, thereby reducing the gap between physical implementation and simulation.
@@ 4.1. Neural Networks for DPP Control
The decoder predicts wavefront coefficients from control voltages using a dual-branch architecture: a linear branch to model overall linearity and a sigmoid-activated branch to capture residual nonlinearity. This residual learning strategy builds upon the linear model’s approximate estimation (Rajaeipour et al., 2021), requiring only nonlinear refinements. The input is extended to $[V,V^{2}]$, to facilitate pattern recognition. Conversely, the encoder, which predicts the desired control voltages from wavefront coefficients, is modeled as a non-linear mapping using a simple two-layer multilayer perceptron with a final sigmoid activation to constrain the output within the physical voltage limits. Voltage normalization by $V_{max}$ ($\tilde{V}=V/V_{max}$) is applied to ensures stable training.
@@ 4.2. Wavefront Measurement
A coded wavefront sensor (CWFS) (Wang et al., 2019) is used to measure wavefront changes in the DPP. As shown in Fig 8, a $4f$ system relays the DPP’s aperture to the CWFS, with alignment ensured by an auto-collimator. Lenses with focal lengths of $150\,\mathrm{mm}$ and $100\,\mathrm{mm}$ provide a spatial magnification of $-2/3$, effectively projecting the full $10\,\mathrm{mm}$ DPP aperture onto most of the CWFS sensor area ($12.46\times 6.7\,\mathrm{mm}$). Illumination is provided by a collimated $520\,\mathrm{nm}$ laser source. The DPP is mounted on a flip-mount that can rotate $90^{\circ}$, allowing it to be removed from the optical path to provide a collimated wavefront reference. The CWFS then measures the DPP’s relative wavefront changes. From the measured relative wavefront, we fit the Zernike polynomials within the circular aperture region to get the corresponding coefficients.
Figure 8. Wavefront Measurement Setup. (a) The optical path of the Measurement system. A 4F system with a spatial magnification of $-2/3$ relays the DPP device’s aperture plane to the wavefront sensor. (b) The realworld experiment setup. (c) The DPP is mounted on a flip-mount that can rotates 90 degrees to remove it from the optical path.
@@ 4.3. Training of DPP Control Models
We captured 1800 voltage-Zernike coefficient pairs by randomly sampling Zernike coefficients within the device’s operating range of $(-3,3)\mu m$. We used $80\%$ of the data for training and the remaining $20\%$ for testing. For comparison, a linear decoding model was trained alongside our neural-network (NN) model. To address measurement error and prevent overfitting, random noise with a standard devitation of 0.36 was added to the Zernike coefficients for data augmentation. As illustrated in Fig 9, the neural network model achieves significantly lower training and testing loss than the linear model, approaching the measurement limit established by repeated CWFS measurements. On the test set, the linear model’s reconstruction error (RMSE) increases with larger amplitudes, indicating poor performance at higher magnitudes, whereas the neural network maintains a more consistent error across magnitudes.
@@ 5. Image Stacking
While fovea stacking shares conceptual similarities with focus stacking, deep learning-based focus stacking methods are often trained on spatial-invariant, disk-pattern blur kernels, which offer limited improvement over traditional sharpness-based techniques. As a result, we adopt a conventional image fusion pipeline. We identify sharp regions by applying a Laplacian filter, followed by blurring. Each pixel in the fused image is computed as a weighted average across the image stack, with weights determined by sharpness. Further implementation details are provided in the supplement. Experiments demonstrate that, despite its simplicity, this stacking method is effective in producing high-quality, aberration-corrected images. A comparison of fusion methods is provided in the Supplement.
@@ 6. Prototype system and Calibration
Our prototype comprises a DPP, a doublet lens, and an image sensor. As shown in Fig. 12, the lens is housed in a zooming enclosure (Thorlab SM1ZM) to adjust the back focus distance ($d_{sensor}$) for different working distances. A 50 mm achromatic doublet lens (Thorlab AC-254-050-A) was selected for its simplicity and inherent achromatic aberration compensation. The Bayer-pattern RGB sensor (FLIR GS3-U3-41C6C-C) features a pixel size of $5.5\mu m\times 5.5\mu m$ with $2048\times 2048$ resolution. Since the components are assembled using threaded lens tubes, the relative distances between optical elements require further calibration.
We calibrate the prototype system using an image-based method that leverages differentiable optics to optimize calibration parameters by matching captured images with simulated counterparts. This process addresses unmeasured distances within the camera setup: the distance between the DPP and the lens’s front plane ($d_{dpp}$), and the distance between the lens’s back plane and the sensor ($d_{sensor}$). The imaging model also incorporates the distance between the image plane and the DPP ($d_{img}$) and the shift between the image center and the intersection of the object plane with the optical axis ($c_{img}$). Given known screen and sensor dimensions, these parameters are jointly optimized. To ensure robust calibration, we capture images using multiple DPP patterns and employ a translation stage to precisely move the image plane. Due to differences in display characteristics and sensor response, the pixel values of captured and simulated images may differ. Therefore, we use the structural similarity index measure (SSIM) to compare local image structure rather than pixel values. All distances are initialized using rough physical measurements, and the image shift $c_{img}$ is calculate from a homography transform estimated by matching the captured image to the known screen pattern.
@@ 7. Experiments
In this section, we experimentally demonstrate the capabilities of the hardware prototype in three applications: aberration-corrected imaging via fovea stacking at far distances (Sec. 7.1), extended depth-of-field imaging via fovea stacking across various depths (Sec. 7.2), and foveated object tracking during smooth pursuit movements (Sec. 7.3).
Figure 13. Fovea stacking: (a) A single capture suffer from significant off-axis aberration. (b) Focal stacking 5 images partially mitigates the curvature of field issue and thus improve the imaging quality, but doesn’t eliminate the aberrations. (c) Fovea stacking provides superior aberration correction. Mask index is visualized for stacking.
@@ 7.1. Fovea Stacking
For imaging beyond hyperfocal distance, the sensor plane is re-positioned via zoom housing rotation. Five phase patterns are then jointly optimized at distance 60m as described in Sec. 3.4 and used for image capture. For comparison, a focus stack with the same number of images was optimized using only defocus as a variable Zernike polynomial term. As shown in Fig. 13, single images are affected by significant off-axis aberrations (Fig. 13a). Although focus stacking can partially compensate for field curvature, it does not fully correct these aberrations (Fig. 13b). Fovea stacking offers better aberration correction, as demonstrated by substantially improved resolution in fine structures and enhanced text readability in this comparison (Fig. 13c). The stacking use the sharpness-based method described in Sec. 5.
@@ 7.2. Fovea Stacking for Extended Depth of Field
To achieve extended depth of field using fovea stacking (Fig. 15), we sampled three planes evenly in disparity space between 535 and 835 mm, optimizing five phases per plane for a total of 15 images. Four objects were placed at varying depths, with two car models positioned at angles to enhance the depth variation. For a fair comparison with the focus stack method, we also optimized 15 phases, varying only the defocus term in the Zernike polynomials.
Fig. 16 shows that a single capture (Fig. 16a) focused around 652mm suffers from significant off-axis and out-of-focus aberrations. While focus stacks (Fig. 16b) mitigate out-of-focus aberrations, off-axis aberrations still limit resolution in peripheral areas. Fovea stacking (Fig. 16c) corrects for both aberrations, and preserves finer details.

### v2

@@ 3.3.1. ROI tiling
Aberration correction across the field of view can be achieved by dividing the FoV into regions of interest (ROI) and correcting local aberrations independently. To optimize each wavefront for local correction within a ROI at distance $D$, $M$ points are sampled - typically on a grid - within the ROI. The optimization loss, $L_{ROI}$, is defined as the mean RMS spot size across these sampled points:
@@ 3.3.2. Joint optimization
Although full-aberration correction can be achieved by tiling a grid of independently corrected ROIs, this approach does not efficiently exploit the information available across the entire image stack. Therefore, to jointly optimize aberration correction across the full FoV with a limited image budget $N$, we propose to optimize the RMS spot size grid across the stack collectively.
We first sample a grid of $H\times W$ points over the entire FoV at a given depth. Ray-tracing is then performed for each of the $N$ different DPP wavefronts, and the corresponding RMS spot size is calculated, resulting in a stack of RMS spot size grid $\{r_{n}(i,j)\in R^{H\times W}|1\leq n\leq N\}$. Each element $r_{n}(i,j)=r_{n}(P_{ij})$ represents the RMS spot size of the sampled point $P_{ij}$, computed using the n-th DPP wavefront. The overall image quality after fusion depends on the minimal RMS spot size at each location across the stack, i.e. $r_{min}(i,j)=\min_{n}r_{n}(i,j)$. Therefore, to jointly optimize for all DPP patterns, we minimize the grid stacking loss.
(4) $L_{gs}=\frac{1}{H\times W}\sum_{i,j}{r_{min}(i,j)}$
Essentially, the grid stacking loss ensures that for each wavefront, only the area where it performs best across the entire stack will be considered as its active region for optimization. However, optimization may yield some degenerate patterns where does not outperform all other waveforms anywhere in the image, thus having no active region at all. This usually happens when the budget of images increases. To alleviate it, we encourage these DPPs to optimize on the hard regions, using $r_{min}/\bar{r}_{min}$ as a weight:
(5) $L_{hr}=$ $\sum_{n^{\prime}}\frac{1}{H\times W}\sum_{i,j}{r_{n^{\prime}}(i,j)\frac{r_{min}(i,j)}{\bar{r}_{min}}}$
$n^{\prime}\in\{n|\sum_{i,j}\mathds{1}\left(n,n^{*}\right)=0\},$
where $\mathds{1}(\cdot)$ is the identity function, and $n^{*}(i,j)=argmin_{n}r_{n}(i,j)$ is the mask index for the sampled point, indicating which DPP wavefront provides the best image quality for this location. The overall loss for joint optimization is a sum of the two losses:
@@
The joint optimization scheme is validated using the same simulated optical system, as shown in Fig. 6. In the experiment, $32\times 32$ grids are sampled across the full FoV to jointly optimize five wavefront patterns. The resulting optimized PSF grids exhibit compensatory patterns due to the gradients being masked differently by the index $n^{*}$ to optimize each DPP wavefront. As expected, the PSF grid displays radial symmetry, which the dynamic optimization scheme efficiently exploits to cover the entire FoV within a limited imaging budget.
@@ 3.3.3. Efficiency under limited imaging budget
The efficiency of using the imaging budget is analyized in Fig 7. With a limited image count, we optimize DPP patterns and evaluate overall imaging quality using the average RMS spot size of the global PSF across the stack—i.e., the mean $r_{min}$ (blue). ROI tiling requires selecting non-overlapping ROIs to cover the entire FoV; thus, image budgets of $2\times 2$, $3\times 3$, and $4\times 4$ are included for comparison (red). Joint optimization significantly improves image quality with as few as 3-5 images, whereas ROI tiling proves less efficient. Beyond 10 images, further improvements in image quality become marginal.
@@ 4. DPP Control Model
where $W\in R^{K}$ represents the Zernike coefficients, and $V^{2}\in R^{63}$ is the element-wise square of the electrode voltages. The matrix $A\in R^{K\times 63}$ is the influence matrix obtained through calibration. To calculate the voltage corresponding to a desired set of Zernike coefficients $W$, it is therefore modeled by solving a constrained optimization problem such that the voltage is within the physical limit $0,V_{max}]$:
[…]
where $V_{max}=270V$ is the maximum voltage for each electrode. This model is also implemented in the drivers of the DPP.
However, this model has two downsides: first, since the hardware capabilities limited by $V_{max}$, working in Zernike space makes it difficult to predict whether a desired shape is feasible on the device. Second, even if the target shape within the feasible range, we find that the linear model breaks down for larger amplitudes. To address this limitation, we propose a neural network-based control strategy that captures the non-linear behavior of the device and improve accuracy, thereby reducing the gap between physical implementation and simulation. Furthermore, by controlling the shape directly in voltage space, the feasibility of the desired wavefront correction is guaranteed.
@@ 4.1. Neural Networks for DPP Control
The decoder predicts wavefront coefficients from control voltages using a dual-branch architecture: a linear branch to model overall linearity and a sigmoid-activated branch to capture residual nonlinearity. This residual learning strategy builds upon the linear model’s approximate estimation (Rajaeipour et al., 2021), requiring only nonlinear refinements. The input is extended to $V,V^{2}]$, to facilitate pattern recognition. Conversely, the encoder, which predicts the desired control voltages from wavefront coefficients, is modeled as a non-linear mapping using a simple two-layer multilayer perceptron with a final sigmoid activation to constrain the output within the physical voltage limits. Voltage normalization by $V_{max}$ ($\tilde{V}=V/V_{max}$) is applied to ensures stable training.
@@ 4.2. Wavefront Measurement
A coded wavefront sensor (CWFS) (Wang et al., 2019) is used to measure wavefront changes in the DPP. As shown in Fig 9, a $4f$ system relays the DPP’s aperture to the CWFS, with alignment ensured by an auto-collimator. Lenses with focal lengths of $150\,\mathrm{mm}$ and $100\,\mathrm{mm}$ provide a spatial magnification of 2/3-2/3, effectively projecting the full $10\,\mathrm{mm}$ DPP aperture onto most of the CWFS sensor area ($12.46\times 6.7\,\mathrm{mm}$). Illumination is provided by a collimated $520\,\mathrm{nm}$ laser source. The DPP is mounted on a flip-mount that can rotate $90^{\circ}$, allowing it to be removed from the optical path to provide a collimated wavefront reference. The CWFS then measures the DPP’s relative wavefront changes. From the measured relative wavefront, we fit the Zernike polynomials within the circular aperture region to get the corresponding coefficients.
Figure 9. Wavefront Measurement Setup. (a) The optical path of the Measurement system. A 4F system with a spatial magnification of 2/3-2/3 relays the DPP device’s aperture plane to the wavefront sensor. (b) The realworld experiment setup. (c) The DPP is mounted on a flip-mount that can rotates 90 degrees to remove it from the optical path.
@@ 4.3. Training of DPP Control Models
We captured 1800 voltage-Zernike coefficient pairs by randomly sampling Zernike coefficients within the device’s operating range. We used $80\%$ of the data for training and the remaining $20\%$ for testing. More details on dataset collection are provided in the supplement. For comparison, a linear decoding model was trained alongside our neural-network (NN) model. To address measurement error and prevent overfitting, random noise with a standard devitation of 0.36 was added to the Zernike coefficients for data augmentation. As illustrated in Fig 10, the neural network model achieves significantly lower training and testing loss than the linear model, approaching the measurement limit established by repeated CWFS measurements. On the test set, the linear model’s reconstruction error (RMSE) increases with larger amplitudes, indicating poor performance at higher magnitudes, whereas the neural network maintains a more consistent error across magnitudes.
@@ 5. Image Stacking
While Fovea Stacking shares conceptual similarities with Focus Stacking, deep learning-based Focus Stacking methods are often trained on spatial-invariant, disk-pattern blur kernels, which offer limited improvement over traditional sharpness-based techniques. As a result, we adopt a conventional image fusion pipeline. We identify sharp regions by applying a Laplacian filter, followed by blurring. Each pixel in the fused image is computed as a weighted average across the image stack, with weights determined by sharpness. Further implementation details are provided in the supplement. Experiments demonstrate that, despite its simplicity, this stacking method is effective in producing high-quality, aberration-corrected images.
@@ 6. Prototype system and Calibration
Our prototype comprises a DPP (Phaseform Delta 7), a doublet lens, and an image sensor. As shown in Fig. 13, the lens is housed in a zooming enclosure (Thorlab SM1ZM) to adjust the back focus distance ($d_{sensor}$) for different working distances. A 50 mm achromatic doublet lens (Thorlab AC-254-050-A) was selected for its simplicity and inherent achromatic aberration compensation. The Bayer-pattern RGB sensor (FLIR GS3-U3-41C6C-C) features a pixel size of $\mum ×\times 5.5 μ\mum$ with $2048\times 2048$ resolution. The total length of the compact imaging system is 15 cm. Since the components are assembled using threaded lens tubes, the relative distances between optical elements require further calibration.
We calibrate the prototype system using an image-based method that leverages differentiable optics to optimize calibration parameters by matching captured images with simulated counterparts. This process addresses unmeasured distances within the camera setup: the distance between the DPP and the lens’s front plane ($d_{dpp}$), and the distance between the lens’s back plane and the sensor ($d_{sensor}$). The imaging model also incorporates the distance between the image plane and the DPP ($d_{img}$) and the shift between the image center and the intersection of the object plane with the optical axis ($c_{img}$). Given known screen and sensor dimensions, these parameters are jointly optimized. To ensure robust calibration, we capture images using multiple DPP patterns and employ a translation stage to precisely move the image plane. Due to differences in display characteristics and sensor response, the pixel values of captured and simulated images may differ. Therefore, we use the structural similarity index measure (SSIM) to compare local image structure rather than pixel values. All distances are initialized using rough physical measurements, and the image shift $c_{img}$ is calculated from a homography transform estimated by matching the captured image to the known screen pattern.
@@ 7. Experiments
In this section, we experimentally demonstrate the capabilities of the hardware prototype in three applications: aberration-corrected imaging via Fovea Stacking at far distances (Sec. 7.1), extended depth-of-field imaging via Fovea Stacking across various depths (Sec. 7.2), and foveated object tracking during smooth pursuit movements (Sec. 7.3). Throughout the experiments, all images were captured in RAW and demosaiced using bilinear interpolation. We did not correct lens distortion because it remains consistent across images, and the DPP tilting terms were set to zero to prevent pixel shifts.
Figure 14. Fovea Stacking: (a) A single capture suffer from significant off-axis aberration. (b) Focus Stacking 5 images partially mitigates the curvature of field issue and thus improve the imaging quality, but doesn’t eliminate the aberrations. (c) Fovea Stacking provides superior aberration correction. Mask index is visualized for stacking.
@@ 7.1. Fovea Stacking
For imaging beyond the hyperfocal distance, the sensor plane is re-positioned via zoom housing rotation. Five phase patterns are then jointly optimized at distance 60m as described in Sec. 3.3 and used for image capture. For comparison, a focus stack with the same number of images was optimized using only defocus as a variable Zernike polynomial term. As shown in Fig. 14, single images are affected by significant off-axis aberrations (Fig. 14a). Although Focus Stacking can partially compensate for field curvature, it does not fully correct these aberrations (Fig. 14b). Fovea Stacking offers better aberration correction, as demonstrated by substantially improved resolution in fine structures and enhanced text readability in this comparison (Fig. 14c). The stacking use the sharpness-based method described in Sec. 5.
@@ 7.2. Fovea Stacking for Extended Depth of Field
To achieve extended depth of field using Fovea Stacking (Fig. 16), we sampled three planes evenly in disparity space between 535 and 835 mm, optimizing five phases per plane for a total of 15 images. Four objects were placed at varying depths, with two car models positioned at angles to enhance the depth variation. For a fair comparison with the Focus Stacking method, we also optimized 15 phases, varying only the defocus term in the Zernike polynomials.
Fig. 17 shows that a single capture (Fig. 17a) focused around 652 mm suffers from significant off-axis and out-of-focus aberrations. While Focus Stacking (Fig. 17b) mitigates out-of-focus aberrations, off-axis aberrations still limit resolution in peripheral areas. Fovea Stacking (Fig. 17c) corrects for both aberrations, and preserves finer details.